OpenStudy (anonymous):

The volume of V of a right circular cylinder can be found b using the formula V= PiR^2h Where R is the radius and H is the height of the cylinder. If a right circular cylinder has a radius of 3 centimeters and contains 54pi cubic centimeters pf water, how many centimeters deep is the water?

OpenStudy (imstuck):

Using the information that you have, can you set up the equation to solve for height?

OpenStudy (anonymous):

V=54pi3^2h does that make sense?

OpenStudy (imstuck):

Ummm....well, it's incorrect. So let's try it together. The formula is\[V=\pi r ^{2}h\]They give you the radius of 3 and the fact that there is 54 pi cubic centimeters in the cylinder. So if you had to make another guess, what does the 54 pi cubic centimeters represent in our formula?

OpenStudy (imstuck):

BRB...just one sec. Don't go anywhere!

OpenStudy (imstuck):

Ok, back now. Anyway, just in case you are not able to reply or something happened to you on your end just like to me on mine...I'll help set it up for you. The 54 pi cubic centimeters is the volume of water, the "V" in your formula. \[54\pi=\pi (3)^{2}h\]When you simplify that you get\[54\pi = 9\pi h\]To solve this for h you divide by 9pi on both sides.\[\frac{ 54\pi }{ 9\pi }=h\]The Pi's cancel out and when you divide 54 by 9 you get 6 cm. So there is 6 cm of water in the cylinder.

OpenStudy (anonymous):

Thank you so much! That really helped! :)

OpenStudy (anonymous):

no problem

OpenStudy (imstuck):

You are very welcome!

OpenStudy (anonymous):

who's talkin to you??